# KiDS-Legacy: Redshift distributions and their calibration

### v1

@@ 7.1 SOM redshift distributions
Figure 12 presents the redshift distributions estimated for the six tomographic bins using the SOM algorithm, for a range of different analysis choices. The results are largely indistinguishable from one-another, except that there are two sets of lines; those using the prior volume weighting and those without. As shown in Sect. 4.3, this is for two reasons. Firstly the prior volume weight acts to smooth the large-scale structure that is imprinted on the redshift distribution of the calibration sample, leading to smoother estimates of the wide-field $N(z)$. Secondly, as shown in Fig. 6, the prior volume weights are able to introduce systematic shifts in the tomographic redshift distributions of the calibration sample before SOM reweighting. The shifts in Fig. 12 are less significant than the example shown in Fig. 6, however, demonstrating that the SOM reweighting has acted to undo some of the prior weight shift.
@@
Finally, we note the similarity between the redshift distributions estimated on the data and on the simulations. In particular bins two to five are essentially identical as estimated on SKiLLS and on the data. Bin six shows the most significant differences: the simulated redshift distribution has a considerable smoothly decreasing tail extending to redshift two, whereas the data $N(z)$ all truncate fairly abruptly at $z\approx 1.6$. We have not explored the origin of this difference here, but note that such a difference will be enhanced by slight differences in the signal-to-noise and size properties of the most distant sources. We have attempted to remove such differences using our matching process, however this difference may indicate that there is still some residual difference between the data and mock galaxies at high redshift.
@@ 7.2 CC redshift distributions
Similar to the mock analysis in Sect. 6.2.3, we measure the clustering redshifts of the KiDS-Legacy lensing sample, compute the joint, inverse-variance weighted CC estimate, and perform the shift-fitting with the fiducial SOM redshift distributions (see above). The final CC measurements are presented in Fig. 13. In general, these are (in part due to the noise adaptation) very similar to the ensemble average of the MICE2 realisations (Fig. 10).
@@ 7.2.1 CCs fitted with SOM redshifts
As a result, the shift-fit parameters $D_{\rm z}^{\rm SOM}$ follow as similar trend as those of MICE2 (Table 7). One major difference is that there is some additional intrinsic scatter, especially at $z>1.0$, where DESI dominates the joint CC measurements, that our simple $f$-term model cannot fully capture. The best-fit $f$-term is just a small fraction of the total uncertainty of the CC measurements [^16] When including smaller measurement scales, the $f$-term contribution becomes significant. (Fig. 13).
This difference is also reflected in the increased uncertainty of $D_{\rm z}^{\rm SOM}$ in the last three tomographic bins, where it reaches $\sigma(D_{\rm z}^{\rm SOM})=0.091$ in bin six. Since we are limited to the maximum redshift of our DESI sample at $z\approx 1.6$, the consequence is that the mean redshift of bin six is not very well constrained by the clustering redshifts. In the five other bins, the uncertainty is at a similar level as MICE2. The magnitude of the shifts is in general smaller for the first three tomographic bins, reaching a maximum of $0.028$ in bin two, which is clearly in agreement with visible offset between the SOM and the CCs in Fig. 13. In bins four and five, the magnitude of the shift is similar to bins two and three, but in bin six it reaches $D_{\rm z}^{\rm SOM}=-0.145$, preferring a shift of the $N_{\rm SOM}(z)$ at low significance to higher redshifts. Overall, only bins one and five are unbiased according to the shift-fit after taking the uncertainties into account.
Figure 13: Comparison of the fiducial data CCs, SOM redshift distributions (used as fit model, re-normalised to the fitted amplitude of the CCs, and additionally with the estimated SOM bias corrected), and shift-fit posterior median and 68 % confidence interval. The additional grey error bar whiskers indicate the total confidence interval that includes the fitted intrinsic scatter.
@@ 7.2.2 Comparison to SOM bias from SKiLLS
As a final test, we can use fitted shift-parameters $D_{\rm z}^{\rm SOM}$ to test how well the SOM calibration of the SKiLLS simulation is representative of the SOM calibration of the KiDS-Legacy data. Similar to our analysis of the MICE2 data (Sect. 6.2.3) we compare $D_{\rm z}^{\rm SOM}$ to the bias of the median SOM redshifts obtained from SKiLLS (see Table 8 and Fig. 14) and find that they are not identical, but closely follow the same trends. The bias is consistent with zero in bin one, positive for bins two to four, and smoothly transitions to negative values in bins five and six. In general, the biases indicated by the $D_{\rm z}^{\rm SOM}$ are somewhat larger than the SOM biases in SKiLLS. However, when factoring in the uncertainties, the difference $\delta z_{\rm med}-D_{\rm z}^{\rm SOM}$ is consistent with zero in all bins, except for bin 4, which exhibits an approximately $2\sigma$ difference, again driven by the CCs preferring larger measured biases.
Table 8: Different redshift bias estimates per tomographic bin obtained from the KiDS-Legacy data. The values listed here are the fiducial mean and median SOM bias obtained from SKiLLS (Sect. 6.1.1), followed by the shift-fit parameter obtained by fitting the CCs with the SOM redshifts, and finally the difference between the median SOM bias and the shift-fit parameter.
@@
3 $-0.001$ / $\phantom{-}0.006\pm 0.010$ $\phantom{-}0.021\pm 0.006$ $-0.014\pm 0.012$
@@
5 $-0.011$ / $-0.005\pm 0.010$ $-0.018\pm 0.028$ $\phantom{-}0.013\pm 0.030$
6 $-0.054$ / $-0.056\pm 0.011$ $-0.145\pm 0.091$ $\phantom{-}0.089\pm 0.092$
171717 The uncertainties of $\delta z_{\rm mean}$ and $\delta z_{\rm med}$ are identical to the third decimal place after applying the error floor. $D_{\rm z}^{\rm SOM}$ is duplicated from Table 7 for comparison.
Figure 14: Comparison of the shift parameters obtained from the KiDS-Legacy data and the SOM bias obtained from SKiLLS. The short gray and blue lines indicate the bias in the mean and median SOM redshift, the red data points the shift-fit parameters $D_{\rm z}$ when fitting CC measurements with the $N_{\rm SOM}(z)$. The black data points represent the difference between the empirical median SOM bias estimate and the shift-fit parameter.
@@ 8 Discussion
The results presented in the previous sections form the basis for measurements of weak gravitational lensing with the KiDS-Legacy data set. Using extensive mock catalogues, we quantify the precision and accuracy of the redshift calibration of the six tomographic bins used in those measurements. In the following, we highlight the most important aspects and lessons learned from this calibration effort.
We rely heavily on mock catalogues that resemble the KiDS and KiDZ data in many important aspects (colour-redshift relation, photometric noise level, photo-$z$ quality, clustering properties, etc.). This reliance makes it necessary to introduce redundancy in the underlying simulations to test the robustness of the results to the assumptions in the creation of the simulations. The two simulations used in this work are quite different. SKiLLS is based on a semi-analytic galaxy model, a full simulation of KiDS/VIKING images, and a replication of the KiDS photometry and shape measurement pipelines on these synthetic images. It also extends to high redshifts ($z<2.3$) covering the whole redshift range of interest for KiDS. Our MICE2 mocks are more simplistic as we do not implement a full image simulation here, employ a parametric model for photometric noise, and add shape measurement weights in a rather ad-hoc way. Also, MICE2 is limited to $z<1.4$, which compromises its ability to calibrate the highest redshifts probed by KiDS. However, it covers a much larger area than SKiLLS, which makes it useful for clustering redshift analyses.
[…]
The CC methodology with the KiDS-Legacy calibration samples is tested extensively on MICE2 and shown to be unbiased within errors if the true $N(z)$ are being used as a model in the shift fit. The goodness-of-fit is satisfactory when we inflate the errors of the CC measurements and the noise according to what we observe on the data. The necessity for this adaptation highlights a possible shortcoming of the MICE2 simulations that do not seem to replicate the full complexity of systematic effects in the data. Still, with this adaptation, the similarity between the CC measurements on the data and simulations gives us confidence in the applicability of the MICE2 results.
@@
The crucial question is then whether the highly complementary primary and secondary validation methods agree on the data. This is answered positively by Fig. 14. The residual bias of the KiDS SOM $N(z)$ suggested by SKiLLS agrees with the bias suggested by shift-fitting the KiDS SOM $N(z)$ to the KiDS CC measurements. The only exception is bin four, which exhibits a $2\sigma$ shift towards negative values, however such a single shift will not have a significant impact on the cosmological results. Both methods of validation are similarly precise in the first three tomographic bins. In bins five and six, the clustering redshifts still suffer from limited calibration samples and possibly further systematics that affect the increasingly faint target samples, e.g. spurious density variations due to variable depth, seeing, etc.
[…]
@@ 9 Summary
In this paper, we present the redshift calibration of the final KiDS WL data set dubbed KiDS-Legacy and based on KiDS-dr5. We develop a calibration strategy that involves multiple levels of redundancy to ensure that we meet the requirement of an accuracy in the mean redshifts of the tomographic bins used for cosmic shear at the per cent level.
The first level of redundancy is represented by the use of two complementary sets of mock catalogues extracted from two quite different types of simulations, SKiLLS and MICE2. Using a newly developed matching algorithm, we arrive at mock catalogues that are highly realistic and emulate the data – the KiDS WL sources as well as spectroscopic calibration samples – with high fidelity.
@@
We show – in essence – that running the SOM on one simulation can be used to calibrate the WL sources in the other simulation with residual bias $\langle\delta_{z}\rangle\la 0.01$. Given that we estimate the match between the more sophisticated simulation, SKiLLS, and the KiDS data to be at least as good as the match between SKiLLS and MICE2, we are confident that SKiLLS can calibrate the SOM $N(z)$ of KiDS at the same level of accuracy or better. The great similarity of the $N(z)$ of the simulated SKiLLS sources and the $N(z)$ estimated with the SOM on the data further justifies the applicability of this conclusion to the KiDS data set.
@@
Additionally, we show that the clustering redshifts are able to correct for any residual bias in the SOM $N(z)$ on the MICE2 simulation. With this result in mind, we run the clustering redshift technique on the data, shift-fitting the SOM $N(z)$ to the clustering measurements, and finding biases that agree with the purely simulated bias estimates from SKiLLS. This again mirrors the results of MICE2 clustering-$z$ vs. SKiLLS SOM $N(z)$, which further validates the realism of the simulations. Passing this strong consistency test suggests a robust calibration and a successful understanding and correction of systematic errors at the per cent level in terms of the mean redshifts of the tomographic bins.
@@
These results will be used to define the (correlated) priors on the mean redshifts of the tomographic bins in the upcoming KiDS-Legacy cosmic shear analyses. A SKiLLS-based SOM $N(z)$ calibration with a conservative error floor of $\langle\delta_{z}\rangle=0.01$ will constitute the fiducial setup. As an alternative, we will also present a purely empirical, somewhat less constraining setup that takes the clustering redshift results as (uncorrelated) priors that make the cosmological conclusions independent of simulations of the redshift calibration.
@@
With the kind of data used here, we reach statistical uncertainties on the mean redshifts with our SOM implementation of $\sigma(\langle\delta_{z}\rangle)\approx 0.002$, which is right in the range of the requirement for Euclid. This suggests that in terms of methodology and calibration data, we are almost ready to calibrate a stage-IV cosmic shear survey. Certainly, the redshift range has to be extended to $z\sim 2$, but this is well within reach. The real challenge will be to reduce the systematic error floor, conservatively estimated here, by about a factor of five. This will require a set of a few highly realistic, complementary simulations that capture the whole complexity of a future cosmic shear experiment.
@@
The statistical uncertainties on the clustering redshifts shown here are still at least a factor of three larger than those Euclid requirements. With an order of magnitude more area in the WL samples and the full power of upcoming wide-field spectroscopic calibration samples (a glimpse is given here with just $\sim 10^{5}$ DESI EDR galaxies), this factor of three is within reach. The systematic error control will be paramount here as well and similarly achieved with redundancy in the simulations that validate the calibration.

### v2

@@ 7.1 SOM redshift distributions
Figure 12 presents the redshift distributions estimated for the six tomographic bins using the SOM algorithm, for a range of different analysis choices. The results are largely indistinguishable from one-another, except that there are two sets of lines; those using the prior volume weighting and those without. As shown in Sect. 4.3, this is for two reasons. Firstly the prior volume weight acts to smooth the large-scale structure that is imprinted on the redshift distribution of the calibration sample, leading to smoother estimates of the wide-field $N(z)$. Secondly, as shown in Fig. 6, the prior volume weights are able to introduce systematic shifts in the tomographic redshift distributions of the calibration sample before SOM re-weighting. The shifts in Fig. 12 are less significant than the example shown in Fig. 6, however, demonstrating that the SOM re-weighting has acted to undo some of the prior weight shift.
@@
Finally, we note the similarity between the redshift distributions estimated on the data and on the simulations. In particular bins two to five are essentially identical as estimated on SKiLLS and on the data. Bin six shows the most significant differences: the simulated redshift distribution has a considerable smoothly decreasing tail extending to redshift two, whereas the data $N(z)$ all truncate fairly abruptly at $z\approx 1.6$. We have not explored the origin of this difference here, but note that such a difference will be enhanced by slight differences in the signal-to-noise and size properties of the most distant sources. We attempted to remove such differences using our matching process; however, this difference may indicate that there is still some residual difference between the data and mock galaxies at high redshift.
@@ 7.2 CC redshift distributions
Similar to the mock analysis in Sect. 6.2.3, we measured the clustering redshifts of the KiDS-Legacy lensing sample, computed the joint, inverse-variance weighted CC estimate, and performed the shift-fitting with the fiducial SOM redshift distributions (see above). The final CC measurements are presented in Fig. 13. In general, these are (in part due to the noise adaptation) very similar to the ensemble average of the MICE2 realisations (Fig. 10).
@@ 7.2.1 CCs fitted with SOM redshifts
As a result, the shift-fit parameters $D_{\rm z}^{\rm SOM}$ follow as similar trend as those of MICE2 (Table 7). One major difference is that we observed some additional intrinsic scatter, especially at $z>1.0$, where DESI dominates the joint CC measurements, which our simple $f$-term model was unable to fully capture. The best-fit $f$-term is a small fraction of the total uncertainty of the CC measurements [^16] When including smaller measurement scales, the $f$-term contribution becomes significant. (Fig. 13).
This difference was also reflected in the increased uncertainty of $D_{\rm z}^{\rm SOM}$ in the last three tomographic bins, where it reached $\sigma(D_{\rm z}^{\rm SOM})=0.091$ in bin six. Since we are limited to the maximum redshift of our DESI sample at $z\approx 1.6$, the consequence is that the mean redshift of bin six was not very well constrained by the clustering redshifts. In the five other bins, the uncertainty was found to be at a similar level as MICE2. The magnitude of the shifts is in general smaller for the first three tomographic bins, reaching a maximum of $0.028$ in bin two, which is clearly in agreement with visible offset between the SOM and the CCs in Fig. 13. In bins four and five, the magnitude of the shift is similar to bins two and three, but in bin six it reaches $D_{\rm z}^{\rm SOM}=-0.145$, preferring a shift of the $N_{\rm SOM}(z)$ at low significance to higher redshifts. Overall, only bins one and five are unbiased according to the shift-fit after taking the uncertainties into account.
Figure 13: Comparison of the fiducial data $C$s, SOM redshift distributions (used as fit model, re-normalised to the fitted amplitude of the CCs, and additionally with the estimated SOM bias corrected), and shift-fit posterior median and 68 % confidence interval. The additional grey error bar whiskers indicate the total confidence interval that includes the fitted intrinsic scatter.
@@ 7.2.2 Comparison to SOM bias from SKiLLS
As a final test, we used the fitted shift-parameters $D_{\rm z}^{\rm SOM}$ to test how well the SOM calibration of the SKiLLS simulation represented the SOM calibration of the KiDS-Legacy data. Similar to our analysis of the MICE2 data (Sect. 6.2.3) we compared $D_{\rm z}^{\rm SOM}$ to the bias of the median SOM redshifts obtained from SKiLLS (see Table 8 and Fig. 14) and found that they are not identical, but closely follow the same trends. The bias is consistent with zero in bin one, positive for bins two to four, and smoothly transitions to negative values in bins five and six. In general, the biases indicated by the $D_{\rm z}^{\rm SOM}$ are somewhat larger than the SOM biases in SKiLLS. However, when factoring in the uncertainties, the difference $\delta z_{\rm med}-D_{\rm z}^{\rm SOM}$ is consistent with zero in all bins, except for bin four, which exhibits an approximately $2\sigma$ difference, again driven by the CCs preferring larger measured biases.
Table 8: Different redshift bias estimates per tomographic bin obtained from the KiDS-Legacy data.  [^17] The values listed here are the fiducial mean and median SOM bias obtained from SKiLLS (Sect. 6.1.1), followed by the shift-fit parameter obtained by fitting the CCs with the SOM redshifts, and finally the difference between the median SOM bias and the shift-fit parameter. The uncertainties of $\delta z_{\rm mean}$ and $\delta z_{\rm med}$ are identical to the third decimal place after applying the error floor. $D_{\rm z}^{\rm SOM}$ is duplicated from Table 7 for comparison.
@@
2 $\phantom{-}0.013$ / $\phantom{-}0.015\pm 0.010$ $\phantom{-}0.028\pm 0.006$ $0.014\pm 0.011$
3 0.001-0.001 / $\phantom{-}0.006\pm 0.010$ $\phantom{-}0.021\pm 0.006$ $0.014\pm 0.012$
4 $\phantom{-}0.008$ / $\phantom{-}0.005\pm 0.010$ $\phantom{-}0.035\pm 0.010$ $0.030\pm 0.014$
@@
[…]
Figure 14: Comparison of the shift parameters obtained from the KiDS-Legacy data and the SOM bias obtained from SKiLLS. The short grey and blue lines indicate the bias in the mean and median SOM redshift, the red data points the shift-fit parameters $D_{\rm z}$ when fitting CC measurements with the $N_{\rm SOM}(z)$. The black data points represent the difference between the empirical median SOM bias estimate and the shift-fit parameter.
@@ 8 Discussion
The results presented in the previous sections form the basis for measurements of WL with the KiDS-Legacy dataset. Using extensive mock catalogues, we quantified the precision and accuracy of the redshift calibration of the six tomographic bins used in those measurements. In the following, we highlight the most important aspects and lessons learned from this calibration effort.
We relied heavily on mock catalogues that resemble the KiDS and KiDZ data in many important aspects (colour-redshift relation, photometric noise level, photo-$z$ quality, clustering properties, etc.). This reliance makes it necessary to introduce redundancy in the underlying simulations to test the robustness of the results to the assumptions in the creation of the simulations. The two simulations used in this work are quite different. SKiLLS is based on a semi-analytic galaxy model, a full simulation of KiDS/VIKING images, and a replication of the KiDS photometry and shape-measurement pipelines on these synthetic images. It also extends to high redshifts ($z<2.3$) covering the whole redshift range of interest for KiDS. Our MICE2 mocks are more simplistic as we do not implement a full image simulation here, employ a parametric model for photometric noise, and add shape-measurement weights in a rather ad hoc way. Also, MICE2 is limited to $z<1.4$, which compromises its ability to calibrate the highest redshifts probed by KiDS. However, it covers a much larger area than SKiLLS, which makes it useful for clustering redshift analyses.
Most importantly, the two simulations are inherently so different that any agreement of the calibration results between SKiLLS (truncated at $z<1.4$ for comparison) and MICE2 can be regarded as a strong sign of systematic robustness. This is exactly what we observe with the SOM calibration for the calibration samples that are most comparable between the two simulations, meaning the samples without lensfit weights (as those weights are not entirely realistic in MICE2); see the $\Delta\langle\delta_{z}\rangle$ rows of scenarios $\left[{\textbf{J}}\right]$ (vs $\left[{\textbf{H}}\right]$) and $\left[{\textbf{K}}\right]$ (vs $\left[{\textbf{I}}\right]$) in Table 5. While the actual calibration sample used in the cosmological analysis looks slightly different (i.e. using shape weights) and the redshift bias values themselves will be different (e.g. compare $\langle\delta_{z}\rangle$ of scenarios $\left[{\textbf{A}}\right]$ and $\left[{\textbf{E}}\right]$), there is no reason to assume that the robustness is affected by these differences. Hence, we use the reported $\Delta\langle\delta_{z}\rangle$ to motivate a conservative systematic error floor of $\delta_{z}=0.01$ for our SOM results. Hence, the SKiLLS SOM results, verified with the MICE2 SOM runs, yield a primary, simulation-based validation of the SOM $N(z)$ on the KiDS data at the per cent level in terms of the mean redshift, which is exactly what is required for the full, uncompromised cosmological exploitation of cosmic shear with KiDS-Legacy.
The CC methodology with the KiDS-Legacy calibration samples is tested extensively on MICE2 and shown to be unbiased within errors if the true $N(z)$ are used as a model in the shift fit. The goodness of fit is satisfactory when we inflate the errors of the CC measurements and the noise according to what we observe on the data. The necessity for this adaptation highlights a possible shortcoming of the MICE2 simulations that do not seem to replicate the full complexity of systematic effects in the data. Still, with this adaptation, the similarity between the CC measurements on the data and simulations gives us confidence in the applicability of the MICE2 results.
@@
The crucial question is then whether the highly complementary primary and secondary validation methods agree on the data. This is answered positively by Fig. 14. The residual bias of the KiDS SOM $N(z)$ suggested by SKiLLS agrees with the bias suggested by shift-fitting the KiDS SOM $N(z)$ to the KiDS CC measurements. The only exception is bin four, which exhibits a $2\sigma$ shift towards negative values; however, such a single shift will not have a significant impact on the cosmological results. The two methods of validation mentioned are similarly precise in the first three tomographic bins. In bins five and six, the clustering redshifts still suffer from limited calibration samples and possibly further systematics that affect the increasingly faint target samples, for example, spurious density variations due to variable depth and seeing.
The residual biases and their uncertainties can be directly translated into priors on the mean redshifts used in the cosmological inference of KiDS cosmic shear measurements. The discussion above motivates at least two main scenarios, one that relies fully on the SOM $N(z)$ and their calibration with SKiLLS  and a second one using the SOM $N(z)$ but calibrated with the CC measurements instead. It is clear that in the latter case the very loose priors on the mean redshifts of bins five and six would severely compromise the constraining power of these bins. So this CC-calibrated setup would constitute a very conservative approach. Even the tighter, SKiLLS-calibrated priors should still be regarded as conservative because the error floor introduced due to residual differences between the SOM runs on SKiLLS (truncated) and MICE2 is erring on the side of caution. There are very good reasons to believe in the superiority of the SKiLLS results. If we took those at face value, we would end up with priors on the mean redshifts that approach the level of completed stage-IV cosmic shear surveys (see row $\sigma_{\delta z}$ of scenario $\left[{\textbf{A}}\right]$ of Table 5).
@@ 9 Summary
In this paper we have presented the redshift calibration of the final KiDS WL dataset, dubbed KiDS-Legacy and based on KiDS-dr5. We developed a calibration strategy that involves multiple levels of redundancy to ensure that we met the requirement of an accuracy of the mean redshifts of the tomographic bins used for cosmic shear at the percent level.
The first level of redundancy is represented by the use of two complementary sets of mock catalogues extracted from two quite different types of simulations, SKiLLS and MICE2. Using a newly developed matching algorithm, we arrive at mock catalogues that are highly realistic and emulate the data (i.e. the KiDS WL sources as well as spectroscopic calibration samples) with high fidelity.
@@
We show, in essence, that running the SOM on one simulation can be used to calibrate the WL sources in the other simulation with residual bias $\langle\delta_{z}\rangle\la 0.01$. Given that we estimate the match between the more sophisticated simulation, SKiLLS, and the KiDS data to be at least as good as the match between SKiLLS and MICE2, we are confident that SKiLLS can calibrate the SOM $N(z)$ of KiDS at the same level of accuracy or better. The great similarity between the $N(z)$ of the simulated SKiLLS sources and the $N(z)$ estimated with the SOM on the data further justifies the applicability of this conclusion to the KiDS dataset.
@@
Additionally, we show that the clustering redshifts are able to correct for any residual bias in the SOM $N(z)$ on the MICE2 simulation. With this result in mind, we ran the clustering redshift technique on the data, shift-fitting the SOM $N(z)$ to the clustering measurements and found biases that agree with the purely simulated bias estimates from SKiLLS. This again mirrors the results of MICE2 clustering-$z$ vs SKiLLS SOM $N(z)$, which further validates the realism of the simulations. Passing this strong consistency test suggests a robust calibration and a successful understanding and correction of systematic errors at the percent level in terms of the mean redshifts of the tomographic bins.
@@
These results have been used to define the priors on the mean redshifts of the tomographic bins in the cosmic shear analyses of KiDS-Legacy (Wright et al., 2025; stoelzner/etal:2025). A SKiLLS-based SOM $N(z)$ calibration with a conservative error floor constitutes the fiducial setup, taking the form of informative correlated Gaussian priors with means $\delta z=\{-2.6,1.4,-0.2,0.8,-1.1,-5.4\}\times 10^{-2}$ and standard deviation $\sigma_{\delta z}\approx 0.01$. As an alternative, we also present a purely empirical, somewhat less constraining setup that takes the clustering redshift results as (differently correlated) priors that make the cosmological conclusions independent of simulations of the redshift calibration.
@@
With the type of data used here, we reach statistical uncertainties on the mean redshifts with our SOM implementation of $\sigma(\langle\delta_{z}\rangle)\approx 0.002$, which falls within the range of the requirement for Euclid. This suggests that, in terms of methodology and calibration data, we are almost ready to calibrate a stage-IV cosmic shear survey. Certainly, the redshift range has to be extended to $z\sim 2$, but this is well within reach. The real challenge will be to reduce the systematic error floor, conservatively estimated here, by about a factor of 5. This will require a set of a few highly realistic, complementary simulations that capture the whole complexity of a future cosmic shear experiment.
@@
The statistical uncertainties on the clustering redshifts shown here are still at least a factor of 3 greater than those required by Euclid. With an order of magnitude more area in the WL samples and the full power of upcoming wide-field spectroscopic calibration samples (a glimpse is given here with just $\sim 10^{5}$ DESI EDR galaxies), this factor of 3 is within reach. Systematic error control will be paramount here as well, and will be similarly achieved through redundancy in the simulations that validate the calibration.